Table below shows a numeric frequency distribution of number of likes on a Facebook fake
news post. Use this information to answer questions 13 and 14.

Likes Cumulative Frequency
400 < 800 7
800 < 1200 21
1200 < 1600 26
1600 < 2000 29
2000 < 2400 30

Question 13

The relative frequency for class 800 < 1200 is:
(1) 0.47
(2) 0.19
(3) 21
(4) 0.7
(5) 14

Question 14

What is the frequency of class 2000 < 2400?
(1) 0.03
(2) 30
(3) 1
(4) 0.97
(5) None of the above.
1
Expert's answer
2020-05-18T18:28:22-0400


a)frequency of class 800 to 1200 will be 21-7= 14

relative frequency = 14/30 = 0.47 ( Option 1)


b)frequency of class 2000 to 2400 will be 30-29 = 1 (option 3)